OpenStudy (anonymous):

plz help....Which value of a solves the equation? a–3 = 8 A.4 B.1/2 C.2 D.-2

OpenStudy (anonymous):

I mean the equation is\[a ^{-3}=8\]

OpenStudy (solomonzelman):

damn, I know the answer, but....

OpenStudy (anonymous):

i'll tell you lol

OpenStudy (anonymous):

i'll help you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[2^{3} = 8\]

OpenStudy (anonymous):

it is b. not a, c, or d

OpenStudy (solomonzelman):

RULE - \(\Huge\color{blue}{ \sf a^{-3}= \frac{1}{a^3} }\) \(\Huge\color{blue}{ \sf 8= \frac{1}{a^3} }\) both sides times a^3, \(\Huge\color{blue}{ \sf 8a^3= 1}\) divide both sides by 8, \(\Huge\color{blue}{ \sf a^3= \frac{1}{8}}\) can you do it from here?

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

im sure it is 1/2 not 1/8. And besides, 1/8 is not one of the choices right

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

its not one of the choices

OpenStudy (anonymous):

@linda3

OpenStudy (linda3):

so you still need help?

OpenStudy (anonymous):

yup dont get what the other ppl are writing :(

OpenStudy (linda3):

oh ok , let me take a look at it! :)

OpenStudy (anonymous):

do you think the answer is C?

OpenStudy (linda3):

hang on let me check

OpenStudy (anonymous):

ok

OpenStudy (linda3):

Okay so this is where you got confused... when the guy told you that a too the power of 3 = 1/8 he did not finish solving the problem. Here you have to find what a is.

OpenStudy (anonymous):

ok

OpenStudy (linda3):

i will show u in a drawing.

OpenStudy (anonymous):

ok :)

OpenStudy (linda3):